OpenStudy (anonymous):

John bought a new computer for $1,650. He paid a $160 down payment and financed the rest for 1 year at an interest rate of 7%. Find the total interest paid on the given amortized loan assuming that John makes monthly payments. Do not round until the final answer. Then, round to the nearest cent

OpenStudy (ranga):

Loan amount = $1650 - $160 = $1490 First calculate the monthly payment using the formula: M = L[c(1 + c)^n] / [(1 + c)^n - 1] M = Fixed Monthly Payment L = Total Loan Amount = $1490 c = interest rate in decimal PER MONTH where the interest rate is assumed to be compounded monthly = 0.07/12 n = total number of monthly payments = 12 (financed for 1 year) Plug in, calculate monthly payment. To find total payment: multiply monthly payment by 12. Subtract the loan amount $1490 to find the interest paid.